OpenStudy (anonymous):

I'm very new here, but I'm stuck on some math questions that I need to get done, can anyone help me? This is the first one: Determine whether the equation represents a direct variation. If it does, find the constant variation. 2y = 5x+1 (If you need the multiple choice answers I'll post them)

OpenStudy (anonymous):

What are the choices?

OpenStudy (mrnood):

do you know what 'direct variation means?

OpenStudy (anonymous):

Not completley.

OpenStudy (anonymous):

Math is kind of tricky for me, I have a hard time understanding.

OpenStudy (anonymous):

What are the choices?

OpenStudy (anonymous):

Here are my options: A.) Not a direct Variation B.) Direct variation, constant of variation is \[\frac{ 5 }{ 2 }\] C.) Direct variation, constant of variation is \[\frac{ 2 }{ 5 }\] D.) Direct variation, constant of variation is \[1 -\frac{ 2 }{ 5 }\]